OpenStudy (anonymous):

what is the solution of the proportion 5/8 = b/12

jimthompson5910 (jim_thompson5910):

Hint: Cross multiply to go from 5/8 = b/12 to 5*12 = 8*b which turns into 60 = 8b

OpenStudy (anonymous):

i have the correct answer but it is b= 7.5 im not sure how i would get that?

jimthompson5910 (jim_thompson5910):

well if 60 = 8b, then 8b = 60 b = 60/8 b = 30/4 b = 15/2 <--- me personally I would stop here b = 7.5

OpenStudy (anonymous):

thank you so much!!

jimthompson5910 (jim_thompson5910):

yw